The width of a rectangle is three fourths of the length. The perimeter of the rectangle becomes 50 cm when the length and the width are each increased by 2 cm. Find the length and the width.
Answer by simranb(63)

Perimeter=2(l+b)
Let the length be x
then width=3/4x
when length and width are increased by 2 the the perimeter=50
2(x+2+3/4x+2)=50
Solved by pluggable solver: EXPLAIN simplification of an expression
Your Result:


YOUR ANSWER


Sorry, this expression cannot be simplified.

Universal Simplifier and Solver


Done!

then length=12
Width=3/4 of x
= 3/4 * 12
= 9
Hope this helps!